# A computable version of Hall’s Harem Theorem and Geometric von Neumann Conjecture

### v3

@@ Abstract.
We prove a version of Hall’s Harem Theorem where the matching realizes a function with some special properties of its cycles. Using this approach we also prove a new computable version of Hall’s Harem theorem. We apply it to non-amenable computable coarse spaces.
@@ 1. Introduction
Analysis of classical mathematical topics from the point of view of complexity of various types is one of the major trends of modern mathematical logic. Amenability is essentially fruitful from this point of view (see [Kec91], [BK96], [KPT05], [MU16], [CK18], [HPP08], [HKP22] and [HKP]). Our research belongs to computable amenability. This is a topic where computable versions of fundamentals of amenability are studied, see papers of M. Cavaleri [Cav17, Cav18], N. Moryakov [Mor18] and the author (together with A. Ivanov) [DI22b, DI22a]. In the present paper we are particularly interested in von Neumann conjecture: the statement that each non-amenable group contains a subgroup isomorphic to $\mathbb{F}_{2}$.
Let $X$ be a set and let $G$ be a group which acts on $X$ by permutations. Then the $G$-space $(G,X)$ is called amenable if there exists a mean $m:\ell^{\infty}(X)\rightarrow\mathbb{R}$ which is invariant under the action of $G$. In particular, if $X=G$, then the group $G$ is called amenable. This definition is due to to J. von Neumann [Neu29]. In 1938, Tarski proved his famous Tarski Alternative Theorem [Tar38], stating that a $G$-space $X$ is amenable if and only if $X$ does not admit a $G$-paradoxical decomposition. Some computable versions of the Tarski Alternative Theorem have been found by the author together with A. Ivanov in [DI22a]. Roughly speaking effective paradoxical deocmpositions of $G$-spaces of computable transformations were introduced and it was proved that their existence is equivalent to non-amenability.
In [Neu29] J. von Neumann proved that amenable groups cannot contain a subgroup isomorphic to the two generated free group $\mathbb{F}_{2}$. The question whether the converse holds i.e. whether each non-amenable group contains a subgroup isomorphic to $\mathbb{F}_{2}$, became popularized as the von Neumann Conjecture in the late 50s (see e.g. [Day57]). It was answered negatively with the construction of Tarski monster group by Olshanskii in 1980 [Ol’81]. However, in 1999 Whyte [Why99] proved so called the geometric version of this conjecture showing existence of $d$-regular forests in non-amenable metric spaces. The latter result is now called Geometric von Neumann Conjecture. The name corresponds to an easy corollary that so called the wobbling group of the space has a subgroup isomorphic to $\mathbb{F}_{2}$ with a semi-regular action. In 2018 Schneider [Sch18] generalized Whyte’s theorem and the corollary to coarse spaces.
@@
  • a computable version of the Hall Harem Theorem.
The Hall Harem Theorem is a generalization of Hall’s marriage theorem, see [CSC10, Theorem H.4.2.]. We consider it together with geometric von Neumann conjecture just because of the well-known fact that aforementioned classical results can be obtained by applications of this theorem.
@@ 1.1. A computable version of Scheider’s theorem
Theorem 1.1 (Geometric Von Neumann Conjecture, [Why99]).
@@
Since we concentrate on Schneider’s generalization of this theorem [Sch18] we skip thorough explanation of the terminology used by Whyte. Schneider’s theorem [Sch18] concerns non-amenable coarse spaces. We will introduce them to the reader. This would give rough understanding of the topic of Geometric von Neumann Conjecture.
Let us recall some terminology from coarse geometry (according to [Sch18]). For a relation $E\subseteq X\times X$ on a set $X$ and $x\in X,A\subseteq X$, let
@@
Note that metric spaces are examples of coarse spaces: for a given metric space $X$, a coarse space $(X,\mathcal{E})$ is obtained by setting:
:=$\mathcal{E}:=\{E\subseteq X\times X|\sup\{d(x,y)|(x,y)\in E\}\leq\infty\}.$
@@
([Sch18, Theorem 2.2]) Let $d\geq 3$. A coarse space $(X,\mathcal{E})$ of bounded geometry is not amenable if and only if there is $E\in\mathcal{E}$ such that $\Gamma(E)$ is a $d$-regular forest.
@@
In the proof of his result, Schneider takes a symmetric entourage $E$ such that $|E(F)|\geq(d-1)|F|$ for all finite $F\subset X$. Such an entourage exists by the fact that when $(X,\mathcal{E})$ is not amenable, then for every finite $d$ there is a symmetric entourage such that $|E(F)|\geq d|F|$ for each finite $F\subseteq X$ (see discussion before Proposition 2.1 in [Sch18]). Now the bipartite graph $\Gamma:=(X,X,R)$ where $R:=E\setminus\Delta_{X}$, satisfies Hall’s Harem condition for $d-1$:
for all finite subsets $Y\subset X$ the following inequality holds:
|$|\{x\,|\,\exists y\in Y(x,y)\in R\}|\geq(d-1)|Y|$.
It follows by Hall’s harem theorem that there exists an entourage which graph is $d$-regular. Moreover, properties of entourages allows us to make it cycle-free without losing $d$-regularity.
@@
Let $d\geq 3$. Let a coarse space $(\mathbb{N},\mathcal{E})$ of a bounded geometry be non-amenable. Furthermore, assume that there exists a highly computable symmetric $E\in\mathcal{E}$ such that for any finite $F\subseteq\mathbb{N}$ we have $|E[F]|\geq(d+2)|F|$. Then there exists a computable $E^{\prime}\in\mathcal{E}$ such that $\Gamma(E^{\prime})$ is a $d$-regular forest. Moreover, there exist an algorithm which for any $m,n\in\mathbb{N}$ recognizes if $m$ and $n$ are in the same connected component of $\Gamma(E^{\prime})$.
To see that this statement is a computable version of Schneider’s result we remind the reader that as we already mentioned when $(X,\mathcal{E})$ is not amenable, then for every finite $d$ there is a symmetric entourage such that $|E(F)|\geq d|F|$ for every finite $F\subseteq X$.
@@
Wobbling groups have attracted growing attention in recent years [JM13], [JdlS15], [Sch18]. One of motivating points is a possibility of a nice reformulation of Whyte’s result in terms of semi-regular subgroups. Recall that a subgroup $G\leq Sym(X)$ is said to be semi-regular if no non-identity element of $G$ has a fixed point in $X$. Theorem 6.1 in [Why99] states that a uniformly discrete metric space $X$ of uniformly bounded geometry is non-amenable if and only if $\mathsf{F}_{2}$ is isomorphic to a semi-regular subgroup of $\mathcal{W}(X)$. Corollary 2.3 of [Sch18] states the same for coarse spaces. Note that these formulations look like von Neumann conjecture!
The following theorem is a computable version of Corollary 2.3 of [Sch18].
@@
Let the coarse space $(\mathbb{N},\mathcal{E})$ of a bounded geometry be non-amenable. Furthermore, assume that there exists a highly computable symmetric $E\in\mathcal{E}$ such that $|E[F]|\geq 6|F|$ for all finite $F\subset\mathbb{N}$. Then there are two computable permutations $\sigma,\pi\in Sym(\mathbb{N})$ such that $\langle\sigma,\pi\rangle$ is a free semi-regular subgroup of $\mathcal{W}(\mathbb{N},\mathcal{E})$.
@@ 1.2. Hall’s harem theorem
The Schneider’s scheme presented above is based on an application of the Hall Harem theorem. The latter describes a condition which is equivalent to existence of a perfect $(1,k)$-matching of a bipartite graph, for example see Section III.3 in [Bol79], Theorem 2.4.2 in [LP86] or Theorem H.4.2 in [CSC10].
In order to obtain a computable version of Schneider’s theorem some special form of Hall’s harem theorem is required. Assume that the graph is of the form $(\mathbb{N},\mathbb{N},E)$, where $E$ denotes the set of edges between natural numbers. Any matching clearly realizes a $k$ to $1$ function, say $f:\mathbb{N}\rightarrow\mathbb{N}$. The main concern of our paper are properties which can be added to such a function. In particular we are especially interested in fast reaching cycles of $f$. This will be formalized below as the property of controlled sizes of cycles. As a result the matching obtained in this way, is realized by a computable subset of $\mathbb{N}^{2}$ with the additional property that there is an algorithm which decides whether two natural numbers $m$ and $n$ belong to the same connected component of $f$.
The special form of Hall’s harem theorem mentioned in the previous paragraph is the heart of our paper and the proof of it (as we evaluate) is 75 per cent of the text. In Sections 2 and 3 (a short one) we give an elementary introduction to both graph theoretic and computability theoretic issues. In Section 4 we formulate the computable version of Hall’s harem theorem, which is our goal. There we also describe the strategy of the proof. It is worth noting here that the proof is in the standard style of building of computably enumerable sets as it is usually done in computability theory. We worn the reader that the construction is technical and difficult. A kind of priority arguments is used in it. The proof of it is given in Sections 3 - 6. In Section 7 we connect this material with computable amenability.
The preprint [Dud21] contains the material of the present paper. It also contains the simplified version of Hall’s harem theorem as above where all computability elements are ruled out. We should mention that these two versions are incomparable, but complement each other so that naturally form two parts of a research project.
@@
In the final remark of this introduction we mention other investigations concerning Hall’s theorems in the context of computability theory. The topic was initiated in papers of Manaster and Rosenstain [Man72] and Kierstead [Kie83]. The latter paper was our starting point. Hall’s theorems have become quite popular in complexity theory. For example see [PZ80], [IRT78].
From the point of view of reverse mathematics Hall’s theorems are studied in [FJLM22], [Hir90] and [HH22].
@@ 2.1. General preliminaries
To introduce the reader to the subject we recall Hall’s Harem theorem. We begin with some necessary definitions from graph theory. We mostly follow the notation of [CSC10]. A graph $\Gamma=(\mathbf{\Gamma},E)$ is called a bipartite graph if the set of vertices $\mathbf{\Gamma}$ is partitioned into sets $U$ and $V$ in such a way, that the set of edges $E$ is a subset of $U\times V$. We denote such a bipartite graph by $\Gamma=(U,V,E)$.
@@
Let $\Gamma=(U,V,E)$. We will say that an edge $(u,v)$ is incident to vertices $u$ and $v$. In this case we say that $u$ and $v$ are adjacent. When two edges $(u,v),(u^{\prime},v^{\prime})\in E$ have a common incident vertex we say that $(u,v),(u^{\prime},v^{\prime})$ are also adjacent. A sequence $x_{1},x_{2},\ldots,x_{n}$ is called a path if each pair $x_{i},x_{i+1}$ is adjacent, $1\leq i<n$.
Below we will denote the set of vertices $\mathbf{\Gamma}$ by the same letter with the graph as a structure, i.e. $\Gamma$. Given a vertex $x\in\Gamma$ the neighbourhood of $x$ is a set
@@
For subsets $X\subset U$ and $Y\subset V$, we define the neighbourhood $N_{\Gamma}(X)$ of $X$ and the neighbourhood $N_{\Gamma}(Y)$ of $Y$ by
@@
For a given vertex $v$ a star of $v$ is a subgraph $S=(\{v\}\cup N_{\Gamma}(v),E^{\prime})$ of $\Gamma$, with $E^{\prime}=((\{v\}\cup N_{\Gamma}(v))\times(\{v\}\cup N_{\Gamma}(v)))\cap E$. A $(1,k)$-fan is a subset of $E$ consisting of $k$ edges incident to some vertex $u\in U$. We say that $u$ is the root of the fan, and when $(u,v)$ belongs to the fan we call $v$ a leaf of it.
@@
The $(1,k)$-matching $M$ is called left perfect (resp. right perfect) if each vertex from $U$ is a root of a fan from $M$ (resp. each vertex from $V$ belongs to exactly one fan of $M$). The $(1,k)$-matching $M$ is called perfect if it is both left and right perfect.
We often view an $(1,k)$-matching as a bipartite graph $M$ where the fan of $u\in U$ is the $M$-star of $u$, i.e. a graph consisting of the set of all vertices adjacent to $u$ in $M$ and all edges incident to $u$ in $M$. We emphasize that a perfect $(1,k)$-matching from $U$ to $V$ is a set $M\subset E$ satisfying following conditions:
@@
    for each vertex $u\in U$ there exists exactly $k$ vertices $v_{1},\ldots,v_{k}\in V$ such that $(u,v_{1}),\ldots,(u,v_{k})\in M$;
@@
    for all $v\in V$ there is an unique vertex $u\in U$ such that $(u,v)\in M$.
Originally the Hall’s Marriage Theorem (see e.g. [Die97, Theorem 2.1.2], [Bol79, Section III.3], [LP86, Theorem 2.4.2.]) provides us a condition for existence of left perfect $(1,1)$-matching in a finite bipartite graph. We are interested in so called Hall’s Harem theorem, which is a generalization of Hall’s Marriage theorem to the case of perfect $(1,k)$-matchings for the locally finite infinite graphs.
One says that graph $\Gamma=(V,E)$ is locally finite if for all vertices $x\in\Gamma$, the neighbourhood $N(x)$ is finite. Note that if $\Gamma$ is locally finite, then $N(X)$ is finite for any finite $X\subset V$.
@@
[CSC10, Theorem H.4.2.] Let $\Gamma=(U,V,E)$ be a locally finite graph and let $k\in\mathbb{N},\;k\geq 1$. The following conditions are equivalent:
@@
It is a crucial fact that the theorem holds for locally finite infinite graphs. At this point we inform the reader that the statement of the Hall Marriage Theorem does not work for infinite graphs in general (see e.g. [Die97, S.2 Ex.6]). Nevertheless, there are versions of this theorem for graphs of any cardinality [Aha88].

### v4

@@ Abstract.
We prove a computable version of the Hall Harem Theorem where the matching realizes a unary function with some special properties of its cycles. We apply it to non-amenable computable coarse spaces. As a result, we obtain a computable version of the geometric von Neumann conjecture.
@@ 1. Introduction
Analysis of classical mathematical topics from the point of view of complexity of various types is one of the major trends of modern mathematical logic. Amenability is essentially fruitful from this point of view (see [kechrisN], [BK], [KPT], [MU], [CK], [HPP], [HKP1] and [HKP2]). Our research belongs to computable amenability. This is a topic where computable versions of fundamentals of amenability are studied; see papers of M. Cavaleri [Cav1, Cav2], N. Moryakov [mor], and the author (together with A. Ivanov) [DuI, CPD]. In the present paper, we are particularly interested in the von Neumann conjecture: the statement that each non-amenable group contains a subgroup isomorphic to $\mathbb{F}_{2}$.
Let $X$ be a set and let $G$ be a group which acts on $X$ by permutations. Then the $G$-space $(G,X)$ is called amenable if there exists a mean $m:\ell^{\infty}(X)\rightarrow\mathbb{R}$ which is invariant under the action of $G$. In particular, if $X=G$, then the group $G$ is called amenable. This definition is due to to J. von Neumann [Neumann1929]. In 1938, Tarski proved his famous Tarski Alternative Theorem [T1], stating that a $G$-space $X$ is amenable if and only if $X$ does not admit a $G$-paradoxical decomposition. Some computable versions of the Tarski Alternative Theorem have been found by the author together with A. Ivanov in [CPD]. Roughly speaking effective paradoxical decompositions of $G$-spaces of computable transformations were introduced and it was proved that their existence is equivalent to non-amenability.
In [Neumann1929] J. von Neumann proved that amenable groups cannot contain a subgroup isomorphic to the two generated free group $\mathbb{F}_{2}$. The question whether the converse holds i.e. whether each non-amenable group contains a subgroup isomorphic to $\mathbb{F}_{2}$, became popularized as the von Neumann Conjecture in the late 1950-s (see e.g. [day]). It was answered negatively with the construction of the Tarski monster group by Olshanskii in 1980 [ols]. However, in 1999 Whyte [Whyte] proved the geometric version of this conjecture showing existence of $d$-regular forests in non-amenable metric spaces. The latter result is now called the Geometric von Neumann Conjecture. The name corresponds to an easy corollary that so-called the wobbling group of the space has a subgroup isomorphic to $\mathbb{F}_{2}$ with a semi-regular action. In 2018 Schneider [schndr] generalized Whyte’s theorem and the corollary to coarse spaces.
@@
  • a computable version of so called Hall’s harem theorem with controlled sizes of cycles.
The Hall Harem Theorem is a generalization of Hall’s marriage theorem, see [csc, Theorem H.4.2.]. We consider it together with the geometric von Neumann conjecture just because of the well-known fact that Tarski’s alternative theorem can be obtained by an application of Hall’s harem theorem. It turns out that this connection lifts to the level of computability [CPD], and, on the other hand, to the geometric level.
@@ 1.1. A computable version of Scheider’s theorem
Theorem 1.1 (Geometric Von Neumann Conjecture, [Whyte]).
@@
Since we concentrate on Schneider’s generalization of this theorem [schndr] we skip thorough explanation of the terminology used by Whyte. Schneider’s theorem [schndr] concerns non-amenable coarse spaces. We will introduce them to the reader. This would give rough understanding of the topic of the Geometric von Neumann Conjecture.
Let us recall some terminology from coarse geometry (according to [schndr]). For a relation $E\subseteq X\times X$ on a set $X$ and $x\in X,A\subseteq X$, let
@@
Note that metric spaces are examples of coarse spaces: for a given metric space $(X,\rho)$, a coarse space $(X,\mathcal{E})$ is obtained by setting:
$\mathcal{E}:=\{E\subseteq X\times X|\sup\{\rho(x,y)|(x,y)\in E\}\leq\infty\}.$
@@
([schndr, Theorem 2.2]) Let $d\in\mathbb{N}$ and $d\geq 3$. A coarse space $(X,\mathcal{E})$ of bounded geometry is not amenable if and only if there is $E\in\mathcal{E}$ such that $\Gamma(E)$ is a $d$-regular forest.
@@
In the proof of his result, Schneider takes a symmetric entourage $E$ such that $|E(F)|\geq(d-1)|F|$ for all finite $F\subset X$. Such an entourage exists by the fact that when $(X,\mathcal{E})$ is not amenable, then for every finite $d$ there is a symmetric entourage such that $|E(F)|\geq d|F|$ for each finite $F\subseteq X$ (see discussion before Proposition 2.1 in [schndr]). Now the bipartite graph $\Gamma:=(X,X,R)$ where $R:=E\setminus\Delta_{X}$, satisfies Hall’s Harem condition for d1d-1:
for all finite subsets $Y\subset X$ the following inequality holds: $|R[Y]|\geq(d-1)|Y|$.
It follows by Hall’s harem theorem that there exists an entourage which graph is $d$-regular. Moreover, the graph can be transformed into a cycle-free member of $\mathcal{E}$ without losing $d$-regularity.
@@
Let $d\geq 3$. Let $(\mathbb{N},\mathcal{E})$ be a non-amenable coarse space of a bounded geometry. Assume that there exists a highly computable symmetric $E\in\mathcal{E}$ such that for every finite $F\subseteq\mathbb{N}$ we have $|E[F]|\geq(d+2)|F|$.
Then there exists a computable $E^{\prime}\in\mathcal{E}$ such that $\Gamma(E^{\prime})$ is a $d$-regular forest. Moreover, there exist an algorithm which for any $m,n\in\mathbb{N}$ recognizes if $m$ and $n$ are in the same connected component of $\Gamma(E^{\prime})$.
To see that this statement is a computable version of Schneider’s result, we remind the reader that as we have already mentioned, when $(X,\mathcal{E})$ is not amenable, then for every finite $d$ there is a symmetric entourage such that $|E(F)|\geq d|F|$ for every finite $F\subseteq X$.
@@
Wobbling groups have attracted growing attention in recent years [Jusch1], [Jusch2], [schndr]. One of motivating points is a possibility of a nice reformulation of Whyte’s result in terms of semi-regular subgroups. Recall that a subgroup $G\leq Sym(X)$ is said to be semi-regular if no non-identity element of $G$ has a fixed point in $X$. Theorem 6.1 in [Whyte] states that a uniformly discrete metric space $X$ of uniformly bounded geometry is non-amenable if and only if $\mathsf{F}_{2}$ is isomorphic to a semi-regular subgroup of $\mathcal{W}(X)$. Corollary 2.3 of [schndr] states the same for coarse spaces. Note that these formulations look like von Neumann’s conjecture!
The following theorem is a computable version of Corollary 2.3 of [schndr].
@@
Let $(\mathbb{N},\mathcal{E})$ be a non-amenable coarse space of a bounded geometry. Assume that there exists a highly computable symmetric $E\in\mathcal{E}$ such that $|E[F]|\geq 6|F|$ for all finite $F\subset\mathbb{N}$.
Then there are two computable permutations $\sigma,\pi\in Sym(\mathbb{N})$ such that $\langle\sigma,\pi\rangle$ is a free semi-regular subgroup of $\mathcal{W}(\mathbb{N},\mathcal{E})$.
@@ 1.2. Hall’s harem theorem
Schneider’s scheme presented above is based on an application of the Hall harem theorem. The latter describes a condition which is equivalent to existence of a perfect $(1,k)$-matching of a bipartite graph; for example, see Section III.3 in [BB], Theorem 2.4.2 in [1986C1] or Theorem H.4.2 in [csc].
In order to obtain a computable version of Schneider’s theorem some special form of Hall’s harem theorem is required. Assume that the graph is of the form $(\mathbb{N},\mathbb{N},E)$, where $E$ denotes the set of edges between natural numbers. Any matching clearly realizes a $k$ to $1$ function, say $f:\mathbb{N}\rightarrow\mathbb{N}$. The main concern of our paper are properties which can be added to such a function. In particular we are especially interested in fast reaching cycles of $f$. This will be formalized below as the property of controlled sizes of cycles. As a result, the matching obtained in this way, is realized by a computable subset of $\mathbb{N}^{2}$ with the additional property that there is an algorithm which decides whether two natural numbers $m$ and $n$ belong to the same connected component of $f$.
The special form of Hall’s harem theorem mentioned in the previous paragraph is the heart of our paper, and the proof of it (as we evaluate) is 75 per cent of the text. In Sections 2 and 3 (a short one) we give an elementary introduction to both graph theoretic and computability theoretic issues. In Section 4 we formulate the computable version of Hall’s harem theorem, which is our goal. There we also describe the strategy of the proof. It is worth noting here that the proof is in the standard style of building of computably enumerable sets, as it is usually done in computability theory. We worn the reader that the construction is technical and difficult. A kind of priority arguments is used in it. The proof of it is given in Sections 5 - 8. In Section 9 we connect this material with computable amenability.
The preprint [HallControlled] contains a version of Hall’s harem theorem with controlled sizes of cycles, where all computability elements are ruled out. Although the proof given in it follows Sections 5 - 8 below, but it is slightly easier. We should mention that these two versions are incomparable. They complement each other and naturally form two parts of a research project. The paper [HallControlled] can be used as a helpful companion of the present paper. Further details concerning these issues are given in Sections 2 and 4 below.
@@
In the final remark of this introduction we mention other investigations concerning Hall’s theorems in the context of computability theory. The topic was initiated in papers of Manaster and Rosenstain [MR] and Kierstead [hak]. The latter paper was our starting point. Hall’s theorems have become quite popular in complexity theory. For example see [NPmat], [SMp].
From the point of view of reverse mathematics Hall’s theorems are studied in [CSM], [MTRM] and [RMMP].
@@ 2.1. General preliminaries
To introduce the reader to the subject we recall Hall’s Harem theorem. We begin with some necessary definitions from graph theory. We mostly follow the notation of [csc]. A graph $\Gamma=(\mathbf{\Gamma},E)$ is called a bipartite graph if the set of vertices $\mathbf{\Gamma}$ is partitioned into sets $U$ and $V$ in such a way, that the set of edges $E$ is a subset of $U\times V$. We denote such a bipartite graph by $\Gamma=(U,V,E)$.
@@
Let $\Gamma=(U,V,E)$. We will say that an edge $(u,v)$ is incident to vertices $u$ and $v$. In this case we say that $u$ and $v$ are adjacent. When two edges $(u,v),(u^{\prime},v^{\prime})\in E$ have a common incident vertex we say that $(u,v),(u^{\prime},v^{\prime})$ are also adjacent. A sequence $x_{1},x_{2},\ldots,x_{n}$ is called a path, if each pair $x_{i},x_{i+1}$ is adjacent, $1\leq i<n$.
Below we will denote the set of vertices $\mathbf{\Gamma}$ by the same letter with the graph as a structure, i.e. $\Gamma$. Given a vertex $x\in\Gamma$, the neighborhood of $x$ is a set
@@
For subsets $X\subset U$ and $Y\subset V$, we define the neighborhood $N_{\Gamma}(X)$ of $X$ and the neighborhood $N_{\Gamma}(Y)$ of $Y$ by
@@
Definition 2.1.
A subset $X$ of $U$ (resp. of $V$) is called connected if for all $x,x^{\prime}\in X$ there exist a path $x=p_{0},p_{1},\ldots,p_{k}=x^{\prime}$ in $\Gamma$ such that $p_{i}\in X\cup N_{\Gamma}(X)$ for all $i\leq k$.
Note here that this definition concerns only bipartite graphs. This notion belongs to Kierstead, [hak]. In the case of connected ordinary graphs we take standard definitions.
For a given vertex $v$, the star of $v$ is the subgraph $S=(\{v\}\cup N_{\Gamma}(v),E^{\prime})$ of $\Gamma$, with
$E^{\prime}=((\{v\}\cup N_{\Gamma}(v))\times(\{v\}\cup N_{\Gamma}(v)))\cap E.$
A $(1,k)$-fan is a subset of $E$ consisting of $k$ edges incident to some vertex $u\in U$. We say that $u$ is the root of the fan, and when $(u,v)$ belongs to the fan we call $v$ a leaf of it.
@@
  • The $(1,k)$-matching $M$ is called left perfect (resp. right perfect) if each vertex from $U$ is a root of a fan from $M$ (resp. each vertex from $V$ belongs to exactly one fan of $M$).
  • The $(1,k)$-matching $M$ is called perfect if it is both left and right perfect.
We often view an $(1,k)$-matching as a bipartite graph $M$ where the fan of $u\in U$ is the $M$-star of $u$, i.e. the star of $u$ in the subgraph $M$. We emphasize that a perfect $(1,k)$-matching from $U$ to $V$ is a set $M\subset E$ satisfying following conditions:
@@
    for each vertex $u\in U$ there exists exactly $k$ vertices $v_{1},\ldots,v_{k}\in V$ such that
    $(u,v_{1}),\ldots,(u,v_{k})\in M$;
@@
    for all $v\in V$ there is a unique vertex $u\in U$ such that $(u,v)\in M$.
Originally Hall’s Marriage Theorem (see e.g. [diestel, Theorem 2.1.2], [BB, Section III.3], [1986C1, Theorem 2.4.2.]) provides us a condition for existence of left perfect $(1,1)$-matching in a finite bipartite graph. We are interested in so called Hall’s Harem theorem, which is a generalization of Hall’s Marriage theorem to the case of perfect $(1,k)$-matchings for the locally finite infinite graphs.
One says that graph $\Gamma=(V,E)$ is locally finite, if for all vertices $x\in\Gamma$, the neighborhood $N(x)$ is finite. Note that if $\Gamma$ is locally finite, then $N(X)$ is finite for any finite $X\subset V$.
@@
[csc, Theorem H.4.2.] Let $\Gamma=(U,V,E)$ be a locally finite graph and let $k\in\mathbb{N},\;k\geq 1$. The following conditions are equivalent:
@@
It is a crucial fact that the theorem holds for locally finite infinite graphs. At this point we inform the reader that the statement of the Hall Marriage Theorem does not work for infinite graphs in general (see e.g. [diestel, S.2 Ex.6]). Nevertheless, there are versions of this theorem for graphs of any cardinality [ahr].